OpenStudy (anonymous):

2 STEP EQUATION 14 - 9x = -8(10 + x)

OpenStudy (anonymous):

We'll do this in much the same way as we solved your last question: 14 - 9x = -8(10 + x) First, we'll distribute the -8 14 - 9x = (-8)(10) + (-8)(x) 14 - 9x = (-80) + (-8x) 14 - 9x = -80 - 8x Now, let's get all the x's on one side and all the regular numbers on the other. We'll add 9x + 80 to both sides of the equation: 14 - 9x = -80 - 8x +80 +9x +80 +9x 94= 1x So x = 94.

OpenStudy (anonymous):

I didn't understand it the last time so thank you!